Representation theory around the channel

The first international workshop on representation theory of groups, finite dimensional algebras and related topics will take place at the

University of Picardie in Amiens, FRANCE

Friday 31 of August to Saturday 1 September 2001

The aim of this workshop is to bring together researchers working in representation theory of global, local and modular group rings, finite dimensional algebras and homological questions such as Hochschild cohomology and to support exchange between these concepts. In particular we want to strengthen the interaction between the research groups in Flanders, in Leicester and in Northern France working in these areas.

There will be around 10 main lectures and enough space for discussion before and after the lectures. We try to encourage in particular young researchers of doctoral and post-doctoral level to participate. The main lectures shall be accessible to a broader audience.

The following researchers have already accepted to give a lecture:

Serge Bouc (Université Paris 7/FRANCE)

Claude Cibils (Université Montpellier 2/FRANCE) Eddy Godelle (Université de Picardie/FRANCE) Eric Jespers (Vrije Universiteit Brussels/BELGIUM) Bernhard Keller (Université Paris 7/FRANCE) Steffen König (University of Leicester/UNITED KINGDOM) Robert Marsh (University of Leicester/UNITED KINGDOM) Nicole Snashall (University of Leicester/UNITED KINGDOM)

For further information please contact

Alexander Zimmermann Faculté de Mathématiques Université de Picardie 33, rue St Leu 80039 Amiens FRANCE alexander.zimmermann@u-picardie.fr

We acknowledge support from CNRS (LAMFA FRE 2270), Université de Picardie, University of Leicester (Mathematics and Computer Science Department), Government of Flanders (PAI tournesol).